### v2

@@ 4.2.1 DYNAMIC WIRELESS FH SPLIT
Latency: Functional split at lower PHY layer imposes stricter latency requirements on FH transmissions due to data exchange between O-RU and O-DU. This becomes particularly critical for uplink transmissions under option 7.2a/7.2b, where the beamforming function resides in the O-RU, while channel estimation and equalization are performed in the O-DU. This split creates a latency-sensitive control loop among these three functions, as shown in the uplink part of Fig. 7. Specifically, the beamforming operation at the O-RU may rely on outdated channel state information (CSI) due to the high latency of transferring CSI back and forth between O-RU and O-DU, which further results in a trade-off between user throughput and FH data rate [ericsson2023massive].
@@
However, in NTNs, the FH propagation delay increases significantly due to the long separation distance between DU and RU [larsen2018survey]. In [seeram2024feasibility], the authors analyzed the maximum propagation distance between O-RU and O-DU when deployed on different NTN platforms. Specifically, the ideal latency requirement for lower PHY layer FH split options should be within 0.25 ms. In [veisi2025non], the authors evaluated the one-way FH latency between O-RU and O-DU, reporting values ranging from 2.1 to 18 ms when deployed on different NTN platforms. Although the FH latency requirement can be relaxed if disabling features such as CSI reporting or reciprocity-based channel estimation in NTN, meeting FH latency constraints while maintaining these features to enhance data rates remains an important challenge.
@@
Cooperation Scale: Multiple FH split options at the PHY layer facilitate coordination among multiple O-RUs to boost performance via distributed MIMO (DMIMO) or coordinated scheduling. For example, considering an ORAN-based NTN architecture with distributed O-RUs and one centralized O-DU, Option 7.2x split has the potential to enable centralized cooperation subject to stringent FH data rate and latency requirements. On the other hand, Option 2, with relaxed requirements on FH data rate or latency, results in limited cooperation capability since each O-RU implements the functions of the PHY and MAC layers locally. Therefore, the trade-off between FH requirements and cooperation among O-RUs should be investigated in the FH split design.
@@
Unlike terrestrial networks, where optical fiber enables centralized cooperation among O-RUs via the O-DU, wireless FH links in NTNs make it considerably more challenging to achieve such centralized cooperation with comparable performance. In [Abdelsadek2022], the authors identified synchronization, outdated CSI, and signalling overhead issues in proposed LEO-based DMIMO networks, where each LEO satellite is assumed to connect to a central unit through ISL. In [Abdelsadek2023], the authors proposed to provide connectivity for handheld devices via LEO DMIMO networks with perfect FH links. Furthermore, [Riera-Palou2022] discussed both centralized and distributed precoding strategies for LEO cell-free MIMO networks, which require various levels of CSI information sharing. Despite these advances, the optimal FH split configuration that enables efficient cooperative processing among NTN platforms remains an open research challenge.
To deal with the FH split issues above, a potential solution is to develop a dynamic FH split scheme based on the NFV characteristic in ORAN-based NTN. It is noted that existing works related to dynamic FH split mainly focused on the data rate and latency requirements in terrestrial networks [Vajd2022, Datsika2021, Diez2021, Villegas2024, pousa2024road, Girycki2024]. Expanding the existing works, we propose a dynamic FH split solution for NTN as shown in Fig. 6. Specifically, the cross-domain SMO supports both hierarchical and hybrid modes for O-RU and O-DU FH split reconfiguration. In the hierarchical mode, the cross-domain SMO transmits FCAPS information over O1 to the O-DU, which uses the Open FH M-Plane interface to manage the O-RU. In the hybrid mode, the cross-domain SMO provides FCAPS information over the Open FH M-Plane interface directly to the O-RU, or over O1 to the O-DU, and then uses the Open FH M-Plane interface to manage the O-RU. For a typical RAN domain operating in the NTN segment, the RAN domain can be modeled as a graph structure with O-RU, O-DU, and O-CU nodes, where the FH and midhaul links serve as edges between O-RU and O-DU nodes, and between O-DU and O-CU nodes, respectively. It should be noted that dynamic FH split optimization is a delay-tolerant service with a long update period. Therefore, the Non-RT RIC can utilize the O1 and Open FH M-plane interfaces to collect the global information for graph construction. To solve the dynamic FH split optimization problem, the hierarchical deep reinforcement learning (DRL) integrated with the graph neural network (GNN) can be a promising solution. Specifically, constrained DRL can be used to ensure the capacity of each FH link, that the sum capacity of FH links is smaller than that of the midhaul link for an O-DU, and that the latency requirement is met. Based on the FH split solution, the hierarchical DRL can be utilized to further determine the cooperation scale among O-RUs, aiming to enhance the quality of service on cell-edge users.
4.2.2 MULTI‑TIER RICS FOR DISTRIBUTED LEARNING
@@
As discussed above, the O-RAN Alliance has introduced Near-RT and Non-RT RICs to manage network control and optimization across different timescales and application scenarios [openRAN-4]. Using the Non-RT control loop, the Non-RT RIC provides guidance, enrichment information, and management of ML models for the Near-RT RIC.
@@
Latency: Near-RT RIC controls the O-CU/O-DU’s functions through the E2 interface and further controls the O-RU’s functions via the Open FH interfaces in O-DU, where the control signal transmissions can be implemented based on optical fiber in the terrestrial network. However, in the NTN scenario, the deployment of Near-RT RIC and O-CU/O-DU/O-RU may be separated with a long mutual distance on different NTN platforms, leading to a non-negligible propagation delay. The delay from Near-RT RIC to O-DU and from O-DU to O-RU remains an important challenge, which will significantly degrade the intelligent Near-RT optimization performance. Additionally, considering the split Option 7.2x, the scheduling function in the O-DU and the beamforming function in the O-RU are typically required to be executed in real-time timescale (<10<\!10\! ms), and the modulation function in the RU needs to be executed in smaller timescale (<1<\!1\! ms) [openRAN-4]. This fact makes it challenging to employ Near-RT RIC when optimizing such functions with a small timescale in NTN.
@@
Distributed Learning: The introduction of the Near-RT RIC in ORAN facilitates the deployment of distributed learning frameworks such as federated learning (FL) and multi-agent deep reinforcement learning (MADRL). In this context, Near-RT RIC serves as the central server to coordinate the training phase among O-RUs and O-DUs. Specifically, the FL framework aims to reduce communication overheads and preserve data privacy. In each FL communication round, O-RUs/O-DUs upload gradients obtained from local data samples to the Near-RT RIC, which aggregates the gradients and updates the global model using a gradient descent algorithm. The updated global model is then broadcast back to all participating O-RUs and O-DUs [farajzadeh2025federated].
Furthermore, the MADRL algorithms extend traditional DRL to enable the coordination among multiple agents. The centralized training and distributed execution (CTDE) has been regarded as a promising solution to achieve the trade-off between performance and communication overheads [deng2025distributed]. Specifically, within the CTDE framework, each O-RU or O-DU shares information with the Near-RT RIC during the training phase to enhance the coordination and cooperation capability. After convergence, each O-RU and O-DU can perform the action inference based on local observation during execution. However, the significant transmission latency in ORAN-based NTN degrades the convergence rate in distributed learning due to multiple rounds of information sharing, and further leads to the asynchronous challenge due to different transmission latencies from Near-RT RIC to O-RUs/O-DUs.
@@
Few studies have focused on distributed learning optimization considering transmission latency in NTNs. In [mahboob2025transforming], the authors proposed an ML-based prediction algorithm with the Near-RT RIC deployed on the ground. However, a terrestrial Near-RT RIC inevitably introduces longer propagation delays and lacks flexibility in continuously optimizing mobile O-DU and O-RU nodes. These limitations highlight the need for a dynamic and platform-aware placement strategy of Near-RT RIC functions. Such a strategy should be based on the set of participating O-DUs/O-RUs, their channel conditions, and the available computational resources of each platform. This observation further reveals a clear edge–cloud trade-off for multi-tier RIC deployment in NTNs. Specifically, deploying RIC at the NTN platforms can significantly reduce control-loop latency, alleviate feeder-link and backhaul loads, and keep detailed KPMs local. However, this approach is constrained by the limited computing and energy resources available on the platform. In contrast, cloud or ground deployment provides sufficient processing capabilities and a global view, but the time-critical control loops may be degraded due to round-trip delays.
@@
A practical solution is the adoption of a multi-tier RIC architecture, as illustrated in Fig. 6. In this architecture, a real-time RIC (RT-RIC) can be implemented within the O-RU or O-DU to control functions, such as modulation, over even shorter timescales [openRAN-TN-1, NTN-challenges-2]. Time-sensitive xApp inference and short control loops are executed in an onboard Near-RT RIC, while computationally intensive tasks, including heavy model training, policy optimization, and cross-constellation coordination, are delegated to the Non-RT RIC or SMO deployed on the ground. Concretely, UAV platforms with strict SWaP constraints can offload computationally intensive AI tasks to ground controllers, whereas HAPs/LEO will host more complex Near-RT functionalities locally and perform model or policy updates less frequently.
4.2.3 FLEXIBLE FUNCTIONAL ROLE SWITCH
@@
To address the challenges above, a promising direction is to design a novel framework that enables the flexible functional role switch, which allows the NTN platforms to switch functional roles to O-RU/O-DU/O-CU based on real-time network conditions and functional demands, as illustrated in Fig. 6. This approach can effectively enhance the scalability of NTNs, and it is implemented based on NFV and the use of COTS components [pousa2024road, hojeij2023dynamic]. Specifically, assuming a HAP initially operates as an O-RU. In situations involving ground O-CU/O-DU failures, excessive FH latency, or emergency response needs, the HAP can dynamically elevate its functional role to an O-DU or even an O-CU by downloading or upgrading the related software. In the meantime, its original O-RU role can be intelligently offloaded to a neighboring HAP through network reconfiguration mechanisms to ensure the completeness of ORAN components. This role-switching capability enhances network adaptability, latency reduction, and service continuity, especially in dynamic or post-disaster environments where traditional terrestrial network may be compromised. However, implementing the role-switching function involves addressing challenges such as computational constraints and power supply to support the tasks executed by the new functional role. For example, the computational loads and the power consumption for O-CU are typically much heavier than those of a single O-RU. Meanwhile, to ensure seamless functional migration, efficient cooperation among different NTN platforms is necessary, which requires synchronous role configuration and a reliable communication link. Furthermore, after the functional migration, the data exchanges in backhaul, midhaul, or FH will also require real-time routing optimization due to the dynamic NTN topology and dynamic RAN topology.
4.2.4 CROSS-DOMAIN SMO
@@
Cross-domain SMO is essential for network planning, dynamic and efficient resource allocation, proactive management, and anticipation of complex topologies in NTNs. It will help to streamline and speed up the process of deploying ORAN-based NTNs for both the integrators and the service providers. Furthermore, it can also reduce both CapEx and OpEx by minimizing the number of isolated operational support systems (OSSs), network management systems (NMSs), and operational tools from different vendors and suppliers across various domains, whenever feasible. However, the current 3GPP and O-RAN Alliance standards do not explicitly define or implement a cross-domain SMO structure. Specifically, 3GPP focuses on network data analytics functions (NWDAF) for the CN domain, while the O-RAN Alliance emphasizes the RICs for the RAN domain. Although 3GPP has introduced the concept of cross-domain management data analytics (MDA), it primarily addresses the machine learning model lifecycle and the processing of management and network data, which is mainly utilized for network-level and delay-tolerant optimization [3gpp28105].
@@
Considering the aforementioned challenges, this paper proposes a cross-domain SMO layer, as illustrated in Fig. 6. Specifically, the cross-domain SMO relies on multiple data analytics functions (DAFs) operating across the TPN and CN. Representative examples include the NWDAF, Jalapeno, and the open network operating system (ONOS). In addition, it also relies on the RICs of the ORAN-based NTNs through the appropriate northbound interfaces [ORAN-management-1]. It is worth noting that this orchestration can be achieved through an intelligent or end-to-end (E2E) software layer. This layer enables seamless integration and unified management of air interface schemes, ORAN-based NTN configurations, hardware and firmware updates, transport management, and core–cloud resources.

### v3

@@ 4-B.2 Latency
Functional split at lower PHY layer imposes stricter latency requirements on FH transmissions due to data exchange between O-RU and O-DU. This becomes particularly critical for uplink transmissions under Option 7.2a/7.2b, where the beamforming function resides in the O-RU, while channel estimation and equalization are performed in the O-DU. This split creates a latency-sensitive control loop among these three functions, as shown in the uplink part of Fig. 7. Specifically, the beamforming operation at the O-RU may rely on outdated channel state information (CSI) due to the high latency of transferring CSI back and forth between O-RU and O-DU, which further results in a trade-off between user throughput and FH data rate [44].
@@
However, in NTNs, the FH propagation delay increases significantly due to the long separation distance between O-RU and O-DU [70]. In [108], the authors analyzed the maximum propagation distance between O-RU and O-DU when deployed on different NTN platforms. Specifically, the ideal latency requirement for lower PHY layer FH split options should be within 0.25 ms. In [124], the authors evaluated the one-way FH latency between O-RU and O-DU, reporting values ranging from 2.1 to 18 ms when deployed on different NTN platforms. Although the FH latency requirement can be relaxed if disabling features such as CSI reporting or reciprocity-based channel estimation in NTN, meeting FH latency constraints while maintaining these features to enhance data rates remains an important challenge.
4-B.3 Cooperation Scale
@@
Multiple FH split options at the PHY layer facilitate coordination among multiple O-RUs to boost performance via distributed MIMO (DMIMO) or coordinated scheduling. For example, considering an ORAN-based NTN architecture with distributed O-RUs and one centralized O-DU, Option 7.2x split has the potential to enable centralized cooperation subject to stringent FH data rate and latency requirements. On the other hand, Option 2, with relaxed requirements on FH data rate or latency, results in limited cooperation capability since each O-RU implements the functions of the PHY and MAC layers locally. Therefore, the trade-off between FH requirements and cooperation among O-RUs should be investigated in the FH split design.
@@
Unlike TN, the wireless FH links in NTNs make it more challenging to achieve efficient centralized cooperation. In [9], the authors identified synchronization, outdated CSI, and signalling overhead issues in proposed LEO-based DMIMO networks, where each LEO satellite is assumed to connect to a central unit through ISL. In [8], the authors proposed to provide connectivity for handheld devices via LEO DMIMO networks with perfect FH links. Furthermore, [102] discussed both centralized and distributed precoding strategies for LEO cell-free MIMO networks, which require various levels of CSI information sharing. Despite these advances, determining the optimal wireless FH split configuration that enables efficient cooperative processing among NTN platforms still remains an open research problem.
To solve the challenges above, we propose a dynamic wireless FH split scheme that leverages NFV capabilities in ORAN-based NTN to address the FH split challenges. Existing works related to dynamic FH split mainly focused on the data rate and latency requirements in TNs [122, 36, 40, 125, 101, 50]. Building on these works, we extend the dynamic FH split solution for NTN, as shown in Fig. 6. Specifically, the cross-domain SMO supports both hierarchical and hybrid modes for O-RU and O-DU FH split reconfiguration. In the hierarchical mode, the cross-domain SMO transmits FCAPS information over O1 to the O-DU, which uses the Open FH M-Plane interface to manage the O-RU. In the hybrid mode, the cross-domain SMO provides FCAPS information over the Open FH M-Plane interface directly to the O-RU, or over O1 to the O-DU, and then uses the Open FH M-Plane interface to manage the O-RU.
To optimize the dynamic wireless FH split performance, the hierarchical deep reinforcement learning (DRL) integrated with the graph neural network (GNN) can be employed as a promising method. For instance, within the NTN segment, a typical RAN domain can be naturally modeled as a graph structure with O-RU, O-DU, and O-CU as nodes, where the FH and midhaul links serve as edges between O-RU and O-DU nodes, and between O-DU and O-CU nodes, respectively. Within the DRL framework, the nodes and edges features of the graph are treated as the system state, the reward is defined according to the optimization objective function, such as throughput or latency, and the action is the FH split option for each O-RU and O-DU pair. It is worth noting that dynamic wireless FH split optimization is a delay-tolerant service with a relatively long update period. As a result, the Non-RT RIC can utilize the O1 and Open FH M-plane interfaces to collect the global network information for graph construction. To guarantee the sum capacity of FH links does not exceed the midhaul capacity of an O-DU and also meet latency requirements, constrained DRL approaches based on Lagrangian methods and action masking techniques can be employed. Based on the FH split solution, the hierarchical DRL can be utilized to further determine the cooperation scale among O-RUs, aiming to enhance the quality of service for cell-edge users.
4-C ENABLER-2: MULTI‑TIER RIC STRUCTURE
@@
As discussed above, the O-RAN Alliance has introduced Near-RT and Non-RT RICs to manage network control and optimization across different timescales and application scenarios [99]. Using the Non-RT control loop, the Non-RT RIC provides guidance, enrichment information, and management of ML models for the Near-RT RIC. To enhance the performance of RICs in ORAN-based NTN, it is essential to consider the impact of non-negligible latency and the system challenges when facilitating the distributed learning frameworks.
4-C.1 Latency
@@
In the NTN scenario, the Near-RT RIC and O-CU/O-DU/O-RU may be deployed separately on different NTN platforms, leading to non-negligible propagation delays. The delay from Near-RT RIC to O-DU and from O-DU to O-RU remains an important challenge, which will significantly degrade the intelligent Near-RT optimization performance. Meanwhile, considering the split Option 7.2x, the scheduling and beamforming functions in O-DU/O-RU are typically required to be executed in a real-time timescale (<10<\!10\! ms), while the modulation function in O-RU needs to be executed in a smaller timescale (<1<\!1\! ms) [99]. This fact makes it challenging to continue employing Near-RT RIC for optimizing such functions with a small timescale.
4-C.2 Distributed Learning
@@
The introduction of the Near-RT RIC in ORAN facilitates the deployment of distributed learning frameworks such as federated learning (FL) and multi-agent deep reinforcement learning (MADRL). In this context, Near-RT RIC serves as the central server to coordinate the training phase among O-RUs and O-DUs. Specifically, the FL framework aims to reduce communication overheads and preserve data privacy. In each FL communication round, O-RUs/O-DUs upload gradients obtained from local data samples to the Near-RT RIC, which aggregates the gradients and updates the global model using a gradient descent algorithm. The updated global model is then broadcast back to all participating O-RUs and O-DUs [46]. Furthermore, the MADRL algorithms extend traditional DRL to enable the coordination among multiple agents. The centralized training and distributed execution (CTDE) has been regarded as a promising solution to achieve the trade-off between performance and communication overheads [38]. Specifically, within the CTDE framework, each O-RU or O-DU shares information with the Near-RT RIC during the training phase to enhance the coordination and cooperation capability. After convergence, each O-RU and O-DU can perform the action inference based on local observation during execution.
However, in the AI-Native ORAN-based NTN framework, ensuring robust and continuously learning AI models, such as MADRL and FL, inevitably requires addressing several inherent system challenges. These challenges include highly heterogeneous transmission latencies, intermittent connectivity, and non-independent and identically distributed data across different NTN orbits. For example, the transmission latencies during multiple rounds of information exchange will slow down the convergence of distributed learning. Meanwhile, heterogeneous latencies among RICs and O-CU/O-DU/O-RU entities may lead to asynchronous updates, thereby disrupting the continuity of AI models.
@@
These challenges highlight the need for a dynamic and platform-aware RIC placement strategy that considers the involved O-DUs/O-RUs, their channel conditions, and each NTN platform’s available computational resources. This observation further reveals the edge–cloud trade-off that must be carefully considered in multi-tier RIC deployment for NTNs. To be more specific, deploying RICs at NTN may help reduce control-loop latency, alleviate feeder-link and backhaul loads, and keep detailed KPMs local, but it is constrained by the limited computing and energy resources on the NTN platform. In contrast, deploying RICs at the cloud or ground can help provide sufficient processing capability and the global information, but the time-critical control loops may be degraded due to round-trip delays.
@@
To solve the challenges above, a practical solution is the adoption of a multi-tier RIC structure, as illustrated in Fig. 6. In this architecture, a real-time RIC (RT-RIC) can be implemented within the O-RU or O-DU to control functions, such as scheduling and beamforming (<10<\!10\! ms), or the modulation function in even shorter timescales (<1<\!1\! ms) [100, 32]. The Near-RT RIC is mainly responsible for supporting time-sensitive xApp inference and short control loops, such as RAN slicing or handover, on a near-real-time timescale ($10\,\text{ms}-1\,\text{s}$). The computationally intensive tasks, including heavy AI model training, policy optimization, and cross-constellation coordination, are delegated to the Non-RT RIC (>1>\!1\! s) deployed on the ground. Concretely, UAV platforms with strict SWaP constraints can offload computationally intensive AI tasks to ground controllers, whereas HAPs/LEO will host more complex Near-RT functionalities locally and perform model or policy updates less frequently. This multi-tier RIC structure ensures the AI models and data sources are accessible across different layers or timescales for training and execution, which further facilitates the zero-touch automation.
4-D ENABLER-3: FLEXIBLE FUNCTIONAL ROLE SWITCH
@@
To address the challenges above, a promising direction is to design a novel framework that enables the flexible functional role switch, which allows the NTN platforms to switch functional roles to O-RU/O-DU/O-CU based on real-time network conditions and functional demands, as illustrated in Fig. 6. This approach is implemented based on NFV and the use of COTS components, which can effectively enhance the scalability and resilience of NTNs [101, 58]. The role switch can typically be triggered by fault-related events or on-demand transition events. Specifically, the fault-related events include O-CU/O-DU failures, excessive FH latency, congestion, insufficient power supply, etc. In the proposed framework, the cross-domain SMO periodically collects telemetry data related to infrastructure, platform status, and ORAN system operation through the O1 and O2 interfaces [77]. By aggregating and analyzing the telemetry data using AI/ML models such as digital twin [83], these events can be detected or even predicted in advance. For the on-demand transition event, when the communication system transitions to serve a different scenario, such as emergency communication under an earthquake or a tsunami, the role switch may also be triggered to provide specific service and maintain service continuity.
To better illustrate the function of role switch, consider a scenario in which a HAP initially operates as an O-RU, as shown in Fig. 8. When triggered by either fault-related events or mission-transition events, the HAP can dynamically elevate its functional role to an O-DU or even an O-CU by downloading or upgrading the related software. In the meantime, its original O-RU role can be intelligently offloaded to a neighboring HAP through network reconfiguration mechanisms, ensuring the completeness of ORAN components and service continuity.
Figure 8: Process of flexible functional role switch.
@@
The role switch enhances network adaptability and latency reduction, especially in dynamic or post-disaster environments where traditional terrestrial networks may be compromised. However, implementing the role switch involves addressing challenges such as computational constraints and power supply to support the tasks executed by the new functional role. For example, the computational loads and the power consumption for O-CU are typically much heavier than those of a single O-RU. Meanwhile, to ensure a seamless functional role switch, different NTN platforms need to maintain efficient cooperation and synchronous role configuration. Moreover, to prevent unauthorized role elevation or offloading, the security issues of role switch need to be addressed through enhanced authentication techniques and continuous cross-domain SMO monitoring. After the functional migration, data exchanges in backhaul, midhaul, or FH will also require real-time routing optimization to deal with the dynamic NTN topology.
4-E ENABLER-4: CROSS-DOMAIN SMO
@@
Cross-domain SMO plays a key role in ORAN-based NTNs by enabling network planning, dynamic and efficient resource allocation, proactive management, and prediction of complicated topologies in NTNs. Despite its importance, current 3GPP or O-RAN Alliance standards have not yet defined or implemented a cross-domain SMO structure. Specifically, 3GPP primarily focuses on the CN domain through the network data analytics functions (NWDAF), while the O-RAN Alliance focuses on the RAN domain via RICs. Although 3GPP has introduced the concept of cross-domain management data analytics (MDA), it mainly addresses the ML model lifecycle and the processing of management and network data, which is mainly utilized for network-level and delay-tolerant optimization, instead of E2E intelligent orchestration [5].
@@
To address this gap, this paper proposes a cross-domain SMO layer, as illustrated in Fig. 6, to enable E2E intelligent orchestration across RAN, TPN, and CN domains. Specifically, the cross-domain SMO relies on multiple data analytics functions (DAFs) operating across different domains. The representative examples include the NWDAF, Jalapeno, and the open network operating system (ONOS). In addition, it also relies on the RICs of the ORAN-based NTNs through the appropriate northbound interfaces [80]. It is worth noting that this orchestration can be achieved through an intelligent or E2E software layer. This layer will support seamless integration and unified management of air interface schemes, ORAN-based NTN configurations, and the updates of hardware and firmware.
The integration of cross-domain SMO layer can streamline and accelerate the deployment of ORAN-based NTNs for both system integrators and service providers. Moreover, it enables more efficient network operation by reducing both CapEx and OpEx. This reduction is achieved by minimizing the number of isolated operational support systems (OSSs), network management systems (NMSs), and operational tools from different vendors across various domains. To illustrate the functionality of the proposed cross-domain SMO, we take the cross-domain slice adaptation as a representative example. To enhance the slice adaptation, a typical control loop spanning the RAN, TPN, and CN can be implemented based on a hierarchical MADRL framework. Specifically, the Non-RT RIC optimizes user-level resource allocation within each slice, the TPN domain controller optimizes the path selection of TPN for each slice, and the CN domain controller optimizes computing resource allocation for different slice categories. Through coordinated decision making across domains, E2E slice adaptation can be achieved to satisfy heterogeneous quality of service (QoS) requirements [73].